Using pictures and/or diagrams, describe in detail the methodology you would use to determine experimentally the...

  1. Using pictures and/or diagrams, describe in detail the methodology you would use to determine experimentally the coefficients of static and kinetic friction between 2 objects. (Hint: Access SMU Physics Demonstrations in the Prescribed Readings link for the Module) (10 points)

please type the answer, please do not write on the paper and post it
Answer #1

The co-efficient of friction (μ) defined as the ratio of frictional force (F) and normal force (N) i.e. reaction force

So, μ = F / N

There are two types of friction co-efficient one is for static friction and another is for kinetic friction.

When friction acts on an object which is at rest is known as static friction. Due to certain amount of applied force, the object is about to move from the rest, is equal to the static force of friction (Fs).

So, co-efficient of static friction μs = Fs / N

If the externally applied force (F) is equal to the force of kinetic friction, (Fk) then the object moves at a constant velocity, and the coefficient of friction involved is called the coefficient of kinetic friction, (μk)

So, co-efficient of kinatic friction μK = FK / N

Horizontal Surface Method is an experiment to determine the values of friction co-efficient:

Refer Fig. 1, it is clear that Object of mass M has a tendency to move to the right because of F = mg.    Fis resisted by the friction force Fk. If m is heavy enough to slide M to the right at constant velocity, then F = Fk . At a constant velocity of M to the right, we may write:

μ k= (Fk / N ) = (F / w) or, μ k  = mg / Mg ,      or,  μ k = m / M .  

This means that the ratio m/M gives us a measure of μk.